1.Which of these fractions is the same as 1/2

1.2/8
2.3/-8
3.2/6
4.3/6

2..
A small rectangular mat has a perimeter of 3m. Its width is 2/3
of its length. Find its dimensions.

1.1.2m;1m
2.0.3m:0.2m
3.0.9m;06m
4.1/3m;2/3m

3.Moyle drives his car 150 km in 3 hours. Find the unit rate.
[1] Moyle drives 50 km per hour.
[2] Moyle drives 1 km per 50 hours.
[3] Moyle drives 150 km per 3 hours
[4] Moyle drives 30 km per hour.
1
Expert's answer
2020-07-28T18:13:52-0400

1."\frac{3}{6}",

when simplified, i.e. by dividing both the numerator and denominator by 3, the answer is"\frac{1} {2}".


2. 0.9m;0.6m

Perimeter of a rectangle =2(L+W)

3=2(L+"\frac{2}{3}" L)

3=(2L+"\frac{4}{3}" L)

3="\frac{10}{3}" L

L="\frac{9}{10}" m

L=0.9m

W="\frac{2}{3}" ×L="\frac{2}{3}" × 0.9=0.6m

Answer is 0.9m;0.6m.

3.Speed ="\frac{Distance} {Time}" ="\frac{150} {3}" =50 km/hr

Thus, Moyle drives 50km per hour